# ONG: Orthogonal Natural Gradient Descent

### v1

@@ 4.3 ONG and ONG+
In this paper, we present ONG and ONG+, extensions of OGD and OGD+ using natural gradients, respectively. Our changes are in red in algorithm 2: we essentially precondition gradients using the Fisher information matrix both when updating model parameters and when storing task gradients in memory (the span of this set is what future gradients are ultimately projected onto).
@@ 4.4 Theoretical Guarantees
This section focuses on studying the ONG algorithm (aka Algorithm 2 without the blue text) to obtain some theoretical results. Analogous to Lemma 3.1 from the original OGD paper  (Farajtabar et al., 2019), we present a proof that the preconditioned, projection-subtracted gradient $\tilde{\boldsymbol{g}}$ is still a descent direction with respect to the Fisher metric.
As described in section 3.1, in the context of probabilistic models, the Riemannian metric tensor is the Fisher information matrix $\boldsymbol{F}$ in model parameter space. Theorem 1 from  Farajtabar et al. (2019) highlights that in this Riemannian manifold, $\boldsymbol{F}^{-1}\nabla\mathcal{L}(\boldsymbol{w})$ is the steepest descent direction of the loss. Thus, if $\langle\tilde{\boldsymbol{g}},-\boldsymbol{F}^{-1}\nabla_{\boldsymbol{w}}\mathcal{L}(\boldsymbol{w})\rangle\leq 0,\tilde{\boldsymbol{g}}$ will be a valid descent direction w.r.t the metric structure of the manifold as well. The following lemma proves this.
@@
Let $\boldsymbol{F}^{-1}\boldsymbol{g}$ be the natural gradient of loss function $\mathcal{L}(\boldsymbol{w})$ and $S=\{\boldsymbol{v}_{1},\ldots,\boldsymbol{v}_{n}\}$ be the orthogonal basis. Define $\tilde{\boldsymbol{g}}=\boldsymbol{F}^{-1}\boldsymbol{g}-\sum_{i}^{k}{\mathrm{proj}_{\boldsymbol{v}_{i}}(\boldsymbol{F}^{-1}\boldsymbol{g})}$, where $\mathrm{proj}_{\boldsymbol{v}_{i}}(\boldsymbol{F}^{-1}\boldsymbol{g})$} denotes the projection of $\boldsymbol{F}^{-1}\boldsymbol{g}$ onto the $\boldsymbol{v}_{i}$. Then, $\tilde{\boldsymbol{g}}$ is also a descent direction for $\mathcal{L}(\boldsymbol{w})$, with respect to the metric structure of the model parameter space.
Proof.
$\langle-\tilde{\boldsymbol{g}},\,\,\boldsymbol{F}^{-1}\boldsymbol{g}\rangle$ $=\langle-\tilde{\boldsymbol{g}},\tilde{\boldsymbol{g}}+\sum_{i=1}^{k}{\mathrm{proj}_{\boldsymbol{v}_{i}}(F^{-1}\boldsymbol{g})}\rangle$ (6)
Since $\tilde{\boldsymbol{g}}=\boldsymbol{F}^{-1}\boldsymbol{g}-\sum_{i=1}^{k}{\mathrm{proj}_{\boldsymbol{v}_{i}}(\boldsymbol{F}^{-1}\boldsymbol{g})}$ is orthogonal to the space spanned by the vectors in $S$ and each $\mathrm{proj}_{\boldsymbol{v}_{i}}(\boldsymbol{F}^{-1}\boldsymbol{g})$ lies in this space, $\tilde{\boldsymbol{g}}$ will be orthogonal to $\sum_{i=1}^{k}{\mathrm{proj}_{\boldsymbol{v}_{i}}(\boldsymbol{F}^{-1}\boldsymbol{g})}$. Thus, $\langle\tilde{\boldsymbol{g}},\sum_{i=1}^{k}{\mathrm{proj}_{\boldsymbol{v}_{i}}(\boldsymbol{F}^{-1}\boldsymbol{g})}\rangle=0$. Substituting this into Eq. 7, we arrive at $\langle-\tilde{\boldsymbol{g}},\,\,\boldsymbol{F}^{-1}\boldsymbol{g}\rangle=-\|\tilde{\boldsymbol{g}}\|^{2}\leq 0$. Therefore, $\tilde{\boldsymbol{g}}$ is a valid descent direction for $\mathcal{L}(\boldsymbol{w})$ while being perpendicular to $S$. ∎
@@ 4.5 EKFAC Fisher Approximation
Calculating the exact Fisher matrix is computationally expensive, making it far too large to use and invert scaling up to millions of parameters with certain neural network models per. A nice compromise between full natural gradient and simple diagonal scaling is the Eigenvalue-corrected Kronecker Factorization (EKFAC)  (George et al., 2018). EKFAC works by providing a diagonal approximation not in the raw parameter basis, but in the Kronecker-factored eigenbasis, where a diagonal captures most of the curvature information at a very low cost. Each layer’s Fisher block is approximated as $\boldsymbol{F}=\boldsymbol{A}\otimes\boldsymbol{B}$ with the Kronecker factors $\boldsymbol{A}=\boldsymbol{Q}_{\boldsymbol{A}}\boldsymbol{\Lambda_{A}}\boldsymbol{Q_{A}^{\top}}\boldsymbol{B}=\boldsymbol{Q_{B}}\boldsymbol{\Lambda_{B}}\boldsymbol{Q_{B}^{\top}}$ and then tracks the diagonal $\boldsymbol{\Lambda_{A}}$ and $\boldsymbol{\Lambda_{B}}$ in that eigenbasis, so each mini-batch update is just two small eigentransforms plus a diagonal rescaling. We utilize this approximation of the Fisher matrix in our code implementation. Note that to force the Fisher matrix to be PD, we add $\lambda\boldsymbol{I}$ to it in order to force the fisher matrix to be invertible and numerically more stable in training. We adapt the EKFAC algorithm  (Vantanapath and contributors, 2020) in our method, leading to faster and more stable training.
@@ 5.1 Datasets
We evaluate our method on two standard benchmark datasets utilized extensively in continual learning literature: Permuted MNIST and Rotated MNIST.
5.1.1 Permuted MNIST
Permuted MNIST generates $K$ tasks by fixing $K$ independent random permutations of the $784$ input pixels. Each task are the original MNIST images with its pixels reshuffled by one of those permutations, forcing the model to learn to classify under radically different orderings without forgetting prior permutations.
5.1.2 Rotated MNIST
Rotated MNIST applies a fixed rotation angle $\theta_{k}$ to every MNIST digit. At task $k$, all training and test images are rotated by $\theta_{k}$, which forces the model to continually adapt to new orientations while remembering earlier orientations with different angles.
@@ 5.2 Evaluation Criteria
For evaluation purposes, we adapt the same metrics as  Bennani et al. (2020): average accuracy and average forgetting. Let $a_{\mathcal{T},k}$ denote the accuracy of the model on the $\mathcal{T}_{k}$ after being trained on the task $\mathcal{T}$.
5.2.1 Average Accuracy
Average accuracy ($A_{T}$) is the accuracy of the model after its trained on task $\mathcal{T}_{T}$. It is defined as:
@@ 5.3 Setup
We use a base model of a three-layer MLP with 100 hidden units in two layers, with about 115,000 total parameters. Each layer uses RELU activation, and the final output are the 10 logits.
After some experimentation, we settle on 3 epochs per task, a learning rate of 0.001, a memory buffer size of 100 samples per task, and batch size 32. The seeds ranged from 0 to 5. No training augmentations were added, and this could be part of a future follow-up. All training and validation was done on 1 NVIDIA A100.
@@ 5.4.1 Forgetting and Average Validation Accuracy
Here, we look at the average forgetting and average validation accuracy, as described above, of the two baseline methods (OGD and OGD+), as well as our natural gradient extensions to each (ONG and ONG+). Table 1 highlights the average forgetting values, while Table 2 highlights the final average validation accuracy.
Table 1: The average forgetting of all the methods we considered (lower is better).
Dataset Permuted MNIST Rotated MNIST
@@
Table 2: The final validation accuracy, averaged over all tasks, of all the methods we considered.
Dataset Permuted MNIST Rotated MNIST
@@ 6 Conclusion and Outlook
In this paper, we introduce a new continual learning algorithm, ONG, that combines the interference-free subspace projections of OGD with the information geometry preconditioning of the natural gradient. We prove ONG retains the descent-direction guarantees under the Fisher metric, and how to implement it using the EKFAC fisher approximation method. However, interestingly enough ONG performed worse on the tasks of Permuted and Rotated MNIST, particularly on Permuted MNIST. While it was unable to outperform its continual learning counterpart in Euclidean space, these results highlight the challenges in combining Fisher preconditioning with orthogonal-projection-based methods.
Moving forward, an exciting way to improve ONG will be to incorporate more insights from the field of Manifold Learning and Information Geometry. One such idea is the concept of parallel transports. From  Li and Ma (2022), we have the following definition.
@@
$P_{x\to y}:T_{x}\mathcal{M}\;\longrightarrow\;T_{y}\mathcal{M}$
@@
$\bigl{\langle}P_{x\to y}\,\xi,\;P_{x\to y}\,\zeta\bigr{\rangle}_{y}\;=\;\langle\xi,\zeta\rangle_{x}\quad\forall\,\xi,\zeta\in T_{x}\mathcal{M}.$
As  Budninskiy et al. (2018) further highlights, parallel transport provides a powerful way of connecting the geometries of nearby points in manifolds with non-trivial curvature. An exciting extension to ONG would be to utilize parallel transport with a well-chosen curve in order to "transport" all natural gradients into the same tangent space and use that space’s inner product for projections and building the orthonormal basis.
Another exciting avenue for future study is to further explore the parameterization-invariance property. This was a major reason we chose natural gradients in this project, so future work studying which model reparameterizations ONG and similar natural-gradient-based methods remain robust to will be an interesting area of research.
Finally, Permuted MNIST and Rotated MNIST used in our experiments were toy datasets, and we believe it is important to analyze the various continual learning approaches with datasets consisting of more diverse, uncorrelated tasks. We hypothesize that the benefits of natural gradient-based methods will be more pronounced in such challenging setups.

### v2

@@ 4.3 ONG and ONG+
In this paper, we present ONG and ONG+, extensions of OGD and OGD+ using natural gradients, respectively. Our changes are in red in Algorithm 2: we precondition gradients using the Fisher information matrix both when updating model parameters and when storing task gradients in memory (the span of this set is what future gradients are ultimately projected onto).
@@ 4.4 Theoretical Guarantees
This section focuses on studying the ONG algorithm (aka Algorithm 2 without the blue text) to obtain some theoretical results. Analogous to Lemma 3.1 from the original OGD paper Farajtabar et al. (2019), we present a proof that the preconditioned, projection-subtracted gradient $\tilde{\boldsymbol{g}}$ is still a descent direction with respect to the Fisher metric.
As described in section 3.1, in the context of probabilistic models, the Riemannian metric tensor is the Fisher information matrix $\boldsymbol{F}$. Theorem 1 from Farajtabar et al. (2019) highlights that in this Riemannian manifold, $\boldsymbol{F}^{-1}\nabla\mathcal{L}(\boldsymbol{w})$ is the steepest descent direction of the loss. Thus, if $\langle\tilde{\boldsymbol{g}},-\boldsymbol{F}^{-1}\nabla_{\boldsymbol{w}}\mathcal{L}(\boldsymbol{w})\rangle\leq 0,\tilde{\boldsymbol{g}}$ will also be a valid descent direction with respect to the metric structure of the manifold. The following lemma proves this.
@@
Let $\boldsymbol{F}^{-1}\boldsymbol{g}$ be the natural gradient of loss function $\mathcal{L}(\boldsymbol{w})$ and $S=\{\boldsymbol{v}_{1},\ldots,\boldsymbol{v}_{n}\}$ be an orthogonal basis. Define $\tilde{\boldsymbol{g}}=\boldsymbol{F}^{-1}\boldsymbol{g}-\sum_{i}^{k}{\mathrm{proj}_{\boldsymbol{v}_{i}}(\boldsymbol{F}^{-1}\boldsymbol{g})}$, where $\mathrm{proj}_{\boldsymbol{v}_{i}}(\boldsymbol{F}^{-1}\boldsymbol{g})$} denotes the projection of $\boldsymbol{F}^{-1}\boldsymbol{g}$ onto the $\boldsymbol{v}_{i}$. Then, $\tilde{\boldsymbol{g}}$ is also a descent direction for $\mathcal{L}(\boldsymbol{w})$, with respect to the metric structure of the model parameter space.
The proof for this lemma can be found in Appendix A.
@@ 4.5 EKFAC Fisher Approximation
Calculating the exact Fisher information matrix (FIM) and its inverse becomes prohibitively expensive as our model size grows. To make computation tractable, we utilize the Eigenvalue-corrected Kronecker Factorization (EKFAC) George et al. (2018). EKFAC works by providing a diagonal approximation to the FIM not in the raw parameter basis, but in the Kronecker-factored eigenbasis, where a diagonal captures most of the curvature information at a very low cost. Each neural net layer’s Fisher block is approximated as $\boldsymbol{F}=\boldsymbol{A}\otimes\boldsymbol{B}$ with the Kronecker factors $\boldsymbol{A}=\boldsymbol{Q}_{\boldsymbol{A}}\boldsymbol{\Lambda_{A}}\boldsymbol{Q_{A}^{\top}}\text{ and }\boldsymbol{B}=\boldsymbol{Q_{B}}\boldsymbol{\Lambda_{B}}\boldsymbol{Q_{B}^{\top}}$. We track just the diagonal matrices $\boldsymbol{\Lambda_{A}}$ and $\boldsymbol{\Lambda_{B}}$ in these eigenbases, so each mini-batch update just becomes two small eigenbasis transformations plus a diagonal rescaling. We utilize this approximation of the Fisher matrix in our code implementation. We adapt this EKFAC algorithm Vantanapath (2020) in our code implementation for approximating the Fisher matrix, leading to faster and more stable training.
@@ 5.1 Datasets
Our preliminary evaluations include two standard benchmarks utilized extensively in continual learning literature: Permuted MNIST and Rotated MNIST. Permuted MNIST generates $K$ tasks by fixing $K$ independent random permutations of the $784$ input pixels. Each task is then the original MNIST images with its pixels reshuffled by one of those permutations, forcing the model to learn to classify under radically different orderings without forgetting prior permutations. Rotated MNIST applies a fixed rotation angle $\theta_{k}$ to every MNIST digit. At task $k$, all training and test images are rotated by $\theta_{k}$, which forces the model to continually adapt to new orientations while remembering earlier orientations with different angles.
@@ 5.2 Evaluation Criteria
For evaluation purposes, we adapt the same metrics as Bennani et al. (2020): average accuracy and average forgetting. Let $a_{T,k}$ denote the accuracy of the model on the $\mathcal{T}_{k}$ after being trained on the task $\mathcal{T}_{T}$. Average accuracy ($A_{T}$) is the accuracy of the model after its trained on task $\mathcal{T}_{T}$. It is defined as:
@@ 5.3 Setup
Details can be found in Appendix B.
@@ 5.4.1 Forgetting and Average Validation Accuracy
Here, we present the average forgetting and average validation accuracy, as described above, of the two baseline methods (OGD and OGD+), as well as our natural gradient extensions to each (ONG and ONG+). Table 1 highlights the average forgetting values and Table 2 highlights the final average validation accuracy.
Table 1: Average forgetting of all the methods we considered (lower is better).
Permuted MNIST Rotated MNIST
@@
Table 2: Final validation accuracy, averaged over all tasks, of all the methods we considered.
Permuted MNIST Rotated MNIST
@@ 6 Conclusion and Outlook
In this work, we proposed and investigated Orthogonal Natural Gradient Descent (ONG), a novel algorithm that incorporates natural gradients with OGD, a continual learning algorithm.
Our theoretical contributions include proving that ONG maintains descent-direction guarantees under the Fisher metric and detailing an efficient implementation using the EKFAC approximation.
Our empirical evaluation, however, yielded a counterintuitive yet insightful result: ONG underperforms its Euclidean counterpart on standard continual learning benchmarks. Rather than a simple failure of the method, we interpret this as a key finding of our study: there exists a fundamental tension between Fisher preconditioning and standard orthogonal projections. This discovery suggests that a naive combination of these methods is not only suboptimal but can be detrimental, indicating that the underlying geometries are not immediately compatible. This finding motivates a clear path for future research focused on reconciling these two geometric perspectives. Our primary hypothesis is that the type of projection we utilize must be chosen to be compatible with the descent directions natural gradient descent uses. Immediate future work will therefore involve a deeper analysis of the training dynamics, the evolution of the task-specific gradient subspaces, and the accuracy of the Fisher estimates. Ultimately, the goal is to develop a geometrically-meaningful way of performing "projection" with natural gradients that preserves previously seen gradient directions and enables high accuracy on older tasks.
To that end, incorporating more advanced concepts from Manifold Learning and Information Geometry presents an exciting direction. The concept of parallel transport (see Appendix C for its definition) is particularly promising, as it provides a principled way to move tangent vectors (i.e., gradients) between different points on the parameter manifold Budninskiy et al. (2018); Li and Ma (2022). By using parallel transport to bring gradients from previous tasks into the same tangent space as the current gradient, it may be possible to define a geometrically consistent inner product and basis for projection.
Additional avenues for research include a more rigorous study of the parameterization-invariance property, a primary motivator for this work, to understand which kind of model reparameterizations our method is robust to. Finally, we posit that the limitations we observed may be partially attributable to the high task correlation in benchmarks like Permuted and Rotated MNIST. We hypothesize that the benefits of geometrically-aware methods will become more pronounced in more challenging and realistic continual learning scenarios involving diverse, uncorrelated task sequences. Future work will extend our empirical validation to these more complex domains.
Appendix A Proof of Lemma 4.1
$\langle-\tilde{\boldsymbol{g}},\,\,\boldsymbol{F}^{-1}\boldsymbol{g}\rangle$ $=\langle-\tilde{\boldsymbol{g}},\tilde{\boldsymbol{g}}+\sum_{i=1}^{k}{\mathrm{proj}_{\boldsymbol{v}_{i}}(\boldsymbol{F}^{-1}\boldsymbol{g})}\rangle$ (7)
Since $\tilde{\boldsymbol{g}}=\boldsymbol{F}^{-1}\boldsymbol{g}-\sum_{i=1}^{k}{\mathrm{proj}_{\boldsymbol{v}_{i}}(\boldsymbol{F}^{-1}\boldsymbol{g})}$ is orthogonal to the space spanned by the vectors in $S$ and each $\mathrm{proj}_{\boldsymbol{v}_{i}}(\boldsymbol{F}^{-1}\boldsymbol{g})$ lies in this space, $\tilde{\boldsymbol{g}}$ will be orthogonal to $\sum_{i=1}^{k}{\mathrm{proj}_{\boldsymbol{v}_{i}}(\boldsymbol{F}^{-1}\boldsymbol{g})}$. Thus, $\langle\tilde{\boldsymbol{g}},\sum_{i=1}^{k}{\mathrm{proj}_{\boldsymbol{v}_{i}}(\boldsymbol{F}^{-1}\boldsymbol{g})}\rangle=0$. Substituting this into Eq. 8, we arrive at $\langle-\tilde{\boldsymbol{g}},\,\,\boldsymbol{F}^{-1}\boldsymbol{g}\rangle=-\|\tilde{\boldsymbol{g}}\|^{2}\leq 0$. Therefore, $\tilde{\boldsymbol{g}}$ is a valid descent direction for $\mathcal{L}(\boldsymbol{w})$ while maintaining orthogonality to $S$. $\blacksquare$
Appendix B Setup
We use the official code implementation of Bennani et al. [2020], which can be obtained here as our starting point for building ONG. We also adapt the EKFAC algorithm provided by Vantanapath [2020] for our efficient Fisher information matrix implementation. Both repositories have a MIT license.
We train a randomly-initialized MLP (using Kaiming/He initialization) with depth $L=3$, widths $n_{0}=784$, $n_{1}=n_{2}=d=100$, $n_{3}=10$, and ReLU activation functions. The total number of parameters is $\sim 115,000$.
After some experimentation, we chose 3 epochs per task, a learning rate of $0.001$, a memory buffer size of $100$ samples per task, and a batch size $B=32$. No training augmentations were added (though this could be part of a future follow-up). All training and validation was done on a single NVIDIA A100.
Appendix C Parallel Transport
@@
$P_{x\to y}:T_{x}\mathcal{M}\;\longrightarrow\;T_{y}\mathcal{M}$ (9)
@@
$\bigl{\langle}P_{x\to y}\,\xi,\;P_{x\to y}\,\zeta\bigr{\rangle}_{y}\;=\;\langle\xi,\zeta\rangle_{x}\quad\forall\,\xi,\zeta\in T_{x}\mathcal{M}.$ (10)